A wild birdseed mix is made by combining 100 lb of millet seed costing $0.50 per pound with sunflower seeds costing $1.10 per pound. How many pounds of sunflower seeds are needed to make a mixture that costs $0.62 per pound?
Answer by TimothyLamb(4379)

x = pounds of sunflower
y = pounds of millet = 100
---
0.50x + 1.10y = 0.62(x + y)
0.50x + 1.10*100 = 0.62(x + 100)
0.50x + 110 = 0.62x + 62
0.62x - 0.50x = 110 - 62
0.12x = 48
x = 48/0.12
x = pounds of sunflower = 400
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---